OpenStudy (anonymous):

simplify the expression 3squared * 3 to the 5

OpenStudy (phi):

you mean 3^2 * 3^5 \[ 3^2 \cdot 3^5 \] one way is expand 3^2 and 3^5 3^2 means 3*3 3^5 means 3*3*3*3*3 so \[ 3^2 \cdot 3^5 = 3 \cdot 3 \cdot 3 \cdot 3\cdot 3\cdot 3\cdot 3 \] now write the answer is "short hand" using 3^ exponent

OpenStudy (anonymous):

so the answer is 3^7 @phi

OpenStudy (phi):

yes

OpenStudy (phi):

if you do that enough times, you realize that you can do a short cut: add the exponents notice that 2+5=7

OpenStudy (anonymous):

thanks @osx @phi

OpenStudy (anonymous):

can u help me with one more @phi

OpenStudy (phi):

did you post it ?

OpenStudy (anonymous):

im gonna right now